# Interpretable Fine-Tuning and Error Indication for Graph Neural Network Surrogate Models

### v3

@@ 3.1.5 Preserving Baseline Performance
Incorporation of the interpretability enhancement does not come for free, and incurs the following costs: (1) there is an additional offline training step due to fine-tuning; and (2) inference of the fine-tuned models is more costly compared to their baseline counterparts, due to the introduction of an additional set of arithmetic operations through the interpretable module attachment. In light of these costs, the advantage is that the fine-tuned models not only provide interpretability, but also maintain the performance of the baseline models in terms of forecasting accuracy and stability characteristics.
To emphasize this point, Fig. 6 displays root-mean-squared errors (RMSE) in single-step predictions for baseline and fine-tuned models as a function of testing set Reynolds numbers. The RMSE is calculated node-wise, is computed separately for each predicted output feature, and is normalized by the inlet velocity to allow for relative comparisons across the range of tested Reynolds numbers. More formally, for the $j$-th output feature (the velocity component), it is defined as the relative error percentage
@@
Since the fine-tuning procedure discussed thus far leverages the same objective function as the baseline – namely, the single-step forecasting error in Eq. 3 – the trends in testing set evaluations for this metric are overall consistent with what was observed during training (see Fig. 3). More specifically, Fig. 6 shows how the baseline is effectively an upper bound for the model error at all evaluation Reynolds numbers. Additionally, the error in fine tuned models approaches the baseline as the node reduction factor $RF$ increases. It should also be noted that the normalized error curves are not flat – after about Re=35,000, the relative errors begin to increase, emphasizing the increase in unsteady flow complexity in this regime. This increase is likely an artifact of the fixed timestep using during training. Alongside higher velocity magnitudes fluctuation in the near-step regions, since the shedding cycle frequencies are more rapid at higher Reynolds numbers, the forecasting task becomes more difficult in a given fixed GNN $\Delta t$ window.
@@
Although the averaged single-step error is valuable when validating the consistency between baseline and fine-tuned model outputs in an a-priori setting, comparison of model performance in an a-posteriori (or rollout [^1] Here, the model is evaluated in an autoregressive context, and therefore accounts for the accumulation of GNN error throughout a series of prediction timesteps (or rollout steps).) setting is required to confirm that the fine-tuning procedure indeed preserves the overall characteristics of the baseline model in a truly predictive context. Such a comparison is given in Fig. 7(a) and (b), which shows rollout predictions for streamwise and vertical velocity components respectively for three different Reynolds numbers. The rollouts are sampled at three fixed spatial (probe) locations in the step cavity region, where the BFS dynamics are most complex: probe 1 is the closest to the step anchor, and reflects the impact of flow separation on the unsteady dynamics, whereas probes 2 and 3 are successively further downstream.
@@
The rollouts ultimately confirm predictive consistency between the fine-tuned GNN and the corresponding baseline – in almost all cases, the predictions provided by the fine-tuned model very closely resemble those of the baseline. Interestingly, the rollouts in Fig. 7 reveal noticeable error buildup in the near-step regions as opposed to further downstream. The degree of error buildup is most prominent at the highest tested Reynolds number, where the onset of instability starts at about 20 rollout steps; this is consistent with the trends in Fig. 6, which implies that accumulation of single-step error is most pronounced at the highest Reynolds numbers.
@@
While the rollout accuracy, particularly near the step, is less than ideal, two important points must be stressed: (a) the GNN timestep $\Delta t$ is prescribed here to be significantly larger than the CFD timestep (as discussed in the previous section); and (b) error accumulation is expected since the objective function used during training and fine-tuning only accounts for the single-step error. Although inclusion of multiple rollout steps during training has been shown in recent work to significantly improve GNN prediction horizons for baseline models [6], such investigations into long-term stability improvements were not explored in this work, as they significantly increase training times and require complex application-specific training schedules. Since the scope of this work is interpretable fine-tuning with respect to a baseline, the objective of Fig. 7 is to highlight how appending the interpretable fine-tuning module to the baseline GNN does not incur significant accuracy penalties in both single-step and rollout contexts. As a result, it can be concluded that the added benefit of masked field interpretability is a meaningful addition to the black-box model.
@@
Figure 7: (a) Rollout predictions for streamwise velocity component at fixed probe locations for baseline GNN (blue squares) and fine-tuned GNN (RF=16, red circles). Solid black line indicates target value. Rows correspond to different initial conditions sourced from the indicated Reynolds numbers; columns correspond to different probe locations at specified x-locations away from step anchor (y-locations of probes coincide with the step height). (b) Same as (a), but for vertical velocity component. (c) Masked field after 1 rollout step for Re=45,589. (d) Masked field after 16 rollout steps for Re=45,589; inlay shows zoom-in on a portion of the masked field.
@@
Figure 7(c) and (d) provide masked field visualizations from the Re=$45,589$ predictions produced by the fine-tuned GNNs, and show the effect of rollout error accumulation on the masked field structures. More specifically, the masked field at the first rollout step in Fig. 7(c) isolates largely coherent structures in the step cavity, as observed in the previous examples (see Fig. 5), whereas the masked field after 16 rollout steps (Fig. 7(d)) includes the emergence of high frequency checkerboard-like patterns that are largely non-physical. These visualizations show a direct correspondence between rollout errors and spatial coherency observed in the identified masked field structures – although more investigation to this end is required, the advantage is that the masked field is generated directly by the model during inference, and it is possible that such high-frequency artifacts can be used to signal oncoming model error accumulation in an a-posteriori context.
@@ 3.2 Error-Identifying GNNs
The results above showcased the interpretability enhancement provided by the fine-tuned model from the angle of feature extraction. More specifically, the structures identified by the fine-tuned GNNs through the masked fields serve as an accessible link between the model architecture and the optimization goal (the mesh-based flow forecast).
The goal of this section is to build on the above by adding more capability to the identified masked fields from the modeling perspective (i.e., providing utility beyond feature extraction). This is accomplished by adding a regularization term, representing a mean-squared error budget, to the objective function during the fine-tuning process. It is shown below how the inclusion of this relatively simple regularization forces the identified nodes in the masked field to not only produce coherent features, but also to coincide with regions that contribute most significantly to the fine-tuned GNN forecasting error, a novel capability not guaranteed in the standard fine-tuning procedure.
@@ 3.2.1 Error Identification via Regularization
A novel error identification strategy is accomplished using a budget regularization that modifies the objective function used during the fine-tuning procedure. This modified objective function adds a regularization term to the standard single-step prediction objective of Eq. 3, and is given by
@@
Figure 8: (Left) Average masked field budgets (Eq. 6) versus testing Reynolds number produced by fine-tuned GNNs using increasing values of $\lambda$. Black curves correspond to streamwise velocity component, and red to vertical velocity component. (Right) Normalized RMSE (Eq. 4) versus testing Reynolds number for baseline GNN (black) and fine-tuned GNNs trained with different levels of budget regularization scaling factor $\lambda$ (markers). All fine-tuned GNNs in figure use $RF=16$. RMSE curves shown for vertical velocity component (trends are consistent for streamwise component).
In Eq. 6, the quantity ${\cal B}$ is the masked field budget, and represents the contribution to the total mean-squared error for the single-step forecast from the nodes identified in the masked field ${\bf M}_{m}^{i}$, where the subscript $m$ denotes dependence on the input flowfield at the $m$-th time step (${\bf X}_{m}$), and the superscript $i$ indexes a node in the graph. As per the definition of the masked field, ${\bf M}_{m}^{i}$ is 1 if the node is identified through the Top-K pooling procedure, and 0 otherwise. Since the regularization term ${\cal L}_{B}$ is cast as the inverse of the budget ${\cal B}$, a model with a small value of ${\cal L}_{B}$ indicates that a large amount of forecasting error is contained in the nodes identified through the masked field. The parameter $\lambda$ in Eq. 5 dictates the influence of the regularization term during the fine-tuning stage, and therefore plays a critical role. Note that setting $\lambda=0$ recovers the non-regularized fine-tuning strategy discussed throughout Sec. 3.1. Through the dependence of ${\cal B}$ on the masked field, the objective is to ensure that – if the regularization strategy succeeds in reducing ${\cal L}_{B}$ – the nodes responsible for a majority of the forecasting error are accessible as an explicit function of the input mesh-based flowfield ${\bf X}_{m}$, resulting in a type of a-posteriori error tagging strategy.
@@ 3.2.2 Impact of Regularization
The effect of the scale factor $\lambda$ on the budget-regularized fine-tuning is shown in Fig. 8 (left), which plots average budget values (single-step prediction error contained in the masked fields) conditioned on both testing set Reynolds number and velocity component for various fine-tuned GNN configurations. The figure illustrates (a) how various degrees of error tagging capability are indeed being provided through the budget regularization procedure in direct proportion to $\lambda$, and (b) the expected relationship between the node reduction factor and budget (higher $RF$ gives lower budget contributions). For example, the standard fine-tuning strategy corresponding to $\lambda=0$ produces masked fields that describe a noticeably lower amount of model error than the regularized ($\lambda>0$) counterparts. Interestingly, the implication is that without budget regularization, the generated masked fields – even for a relatively large node reduction factor of $RF=4$, meaning that $25\%$ of the total nodes are sampled – account for only roughly half of the forecasting error contribution. On the other hand, as $\lambda$ is progressively increased from $10^{-4}$ to $10^{-2}$, the budget values increase across the board. The $\lambda=10^{-2}$ case nearly reaches a saturation point, where a majority of the single-step forecasting error in most cases is described by the identified nodes. The budget curves are almost flat, with some exceptions at the higher tested Reynolds numbers, implying that the regularization strategy learns a robust and interpretable error-identifying mechanism.
@@
Although the increase in $\lambda$ translates to an increase in error-identifying capability in the fine-tuned GNNs, Fig. 8 (right) shows how the regularization incurs an accuracy tradeoff. More specifically, Fig. 8 (right) shows effects of the $\lambda$ parameter on the fine-tuned GNNs in terms of single-step normalized RMSE values (Eq. 4) conditioned on the same testing set Reynolds numbers. When the budget regularization is introduced, the scaling parameter effectively increases the forecasting errors for all observed Reynolds numbers, implying a type of accuracy tradeoff with respect to the additional modeling capability provided by the budget regularization. This is overall consistent with the $\lambda$-scaling process, in that a greater emphasis on minimizing the inverse budget via high $\lambda$ necessitates less optimization focus on the direct forecasting error objective. This attribute is directly apparent in Fig. 8 (right) – as $\lambda\rightarrow 0$, the RMSE curves converge to the standard fine-tuned GNN, which achieves lower forecasting error relative to the baseline GNN used to initiate the fine-tuning procedure.
@@
To further assess the effect of budget regularization in a more practical inference-based setting, Fig. 9 displays forecasting results in terms of masked field budget and predicted velocity versus time step using an initial condition sourced from the Re=35,392 testing set trajectory. The figure compares fine-tuned GNN forecasts in single-step and rollout prediction configurations, with models corresponding to node reduction factors (RFs) of 16 (the adaptive pooling procedure isolates sub-graphs comprised of 16x fewer nodes than the original).
@@
The single-step forecasts largely mimic the trends observed in Fig. 8 (left), in that the regularization imposed during fine-tuning manifests through an increase in error budget during inference with higher $\lambda$. These single-step forecasts, which assume no accumulation of model forecasting error through forward time steps, highlight the control provided by the regularization procedure by the $\lambda$ parameter. More specifically, Fig. 9 indicates diminishing returns in the regularization procedure with respect to $\lambda$: for example, the jump in budget when moving from $\lambda=0$ to $\lambda=10^{-4}$ is on average higher than when moving from $\lambda=10^{-3}$ to $\lambda=10^{-2}$. The rollout trends, which do account for the accumulation of single-step forecasting error through trajectory generation, further point to this aspect of diminishing returns, and directly reveal how the added emphasis on budget regularization through higher $\lambda$ values incurs an accuracy tradeoff. For example, the fine-tuned GNN trained with $\lambda=10^{-2}$ departs from the target solution much earlier than the others, which is consistent with the fact that it places more emphasis on minimizing the inverse budget during fine-tuning – ultimately, the cost of having a more “useful” masked field through the ability to tag regions of high error is a drop in forecasting accuracy and stability. Interestingly, although the correlation between budget and the $\lambda$ parameter is present for the first few rollout steps, this correlation decays once error starts to pollute the predictions. In other words, the budgets in the regularized models eventually tend towards the budgets produced by the non-regularized model. Since the baseline and fine-tuning procedures utilize single-step training objectives, the impact of error accumulation after a single forecasting step is high – as mentioned earlier, these effects can be mitigated with more robust (albiet more expensive) training strategies that account for the error of more than one future time-step in the objective function.
@@
Figure 9: (Top row) Masked field budgets (Eq. 6) versus time step for single-step (left plot) and rollout (right plot) predictions using initial condition from Re=35,392 testing set trajectory. Fine-tuned GNNs have RF=16. (Bottom row) Corresponding velocity field predictions (vertical component) sampled at a single spatial location (probe 2, same as used in middle row of Fig. 7(b)). The $\lambda=-L_{\text{BL}}$ setting (green squares) is discussed in Sec. 3.2.3.
@@
Figure 10: (Left) Masked fields produced by fine-tuned GNNs (RF=16) of increasing $\lambda$ settings for a given input ${\bf X}_{m}$ (input comes from trajectory described in Fig. 9). (Middle) Streamwise component of the forecast error (${\bf X}_{m+1}-\widetilde{\bf X}_{m+1}$) after a single step prediction for each of the corresponding fine-tuned models. (Right) Vertical component of the error.
@@
It should be noted that for small values of $\lambda$ (e.g., $\lambda=10^{-4}$ in Fig. 9), the regularization in Eq. 5 produces error-identifying masked fields at relatively minimal expense in terms of model accuracy with respect to the baseline GNN, thereby offering a unique interpretable modeling advantage from the perspective of error tagging. This advantage is more apparent when considering Fig. 10, which shows masked fields produced by the fine-tuned GNNs used to make the predictions in Fig. 9. The introduction of $\lambda$ incurs a noticeable change to the identified structures in the masked field; more specifically, the non-regularized $\lambda=0$ fine-tuned GNNs isolate structures in the far downstream region, whereas the regularized models isolate more nodes in the near-step vicinity, particularly near the corner at which flow separation is initiated. The difference in masked field structures emphasizes the difference in GNN fine-tuning objectives, in that the identified nodes in the $\lambda>0$ models represent regions in which the predicted mesh-based flowfield $\widetilde{X}_{m+1}$ will contain a majority of the model error. This is evidenced in the corresponding error fields, which showcases the effect of budget regularization on the localization of forecasting error towards the masked regions.
@@
Additionally, the structures in the masked fields from $\lambda=10^{-4}$ to $\lambda=10^{-2}$ in Fig. 10 are largely similar, which allows the user to tune the degree of error identification capability provided by the fine-tuned GNN. For example, the $\lambda=10^{-4}$ model offers (a) comparable performance to the baseline GNN, (b) interpretability through the identification of coherent structures in the masked field, a utility also provided in the non-regularized $\lambda=0$ model, and (c) the ability to attribute the identified regions to expected sources of forecasting error. Although this latter capability is limited to the scope of the prediction horizon of the model (this is apparent through the rollout results in Fig. 9), the inclusion of the budget-based regularization reveals a promising pathway for understanding and isolating sources of model error during inference through the adaptive pooling strategy.

### v4

@@ 3.1.5 Preserving Baseline Performance
Incorporation of the interpretability enhancement does not come for free, and incurs the following costs: (1) there is an additional offline training step; and (2) inference of the interpretable models is more costly compared to their baseline counterparts, due to the introduction of an additional set of arithmetic operations through the interpretable module attachment. In light of these costs, the advantage is that the interpretable models also maintain the performance of the baseline models in terms of forecasting accuracy and stability characteristics.
To emphasize this point, Fig. 6 displays root-mean-squared errors (RMSE) in single-step predictions for baseline and interpretable models as a function of testing set Reynolds numbers. The RMSE is calculated node-wise, is computed separately for each predicted output feature, and is normalized by the inlet velocity to allow for relative comparisons across the range of tested Reynolds numbers. More formally, for the $j$-th output feature (the velocity component), it is defined as the relative error percentage
@@
Since the interpretability enhancement procedure discussed thus far leverages the same objective function as the baseline – namely, the single-step forecasting error in Eq. 3 – the trends in testing set evaluations for this metric are overall consistent with what was observed during training (see Fig. 3). More specifically, Fig. 6 shows how the baseline is effectively an upper bound for the model error at all evaluation Reynolds numbers. Additionally, the error in interpretable models approaches the baseline as the node reduction factor $RF$ increases. It should also be noted that the normalized error curves are not flat – after about Re=35,000, the relative errors begin to increase, emphasizing the increase in unsteady flow complexity in this regime. This increase is likely an artifact of the fixed timestep using during training. Alongside higher velocity magnitudes fluctuation in the near-step regions, since the shedding cycle frequencies are more rapid at higher Reynolds numbers, the forecasting task becomes more difficult in a given fixed GNN $\Delta t$ window.
@@
Although the averaged single-step error is valuable when validating the consistency between baseline and interpretable model outputs in an a-priori setting, comparison of model performance in an a-posteriori (or rollout [^1] Here, the model is evaluated in an autoregressive context, and therefore accounts for the accumulation of GNN error throughout a series of prediction timesteps (or rollout steps).) setting is required to confirm that the interpretability enhancement procedure indeed preserves the overall characteristics of the baseline model in a truly predictive context. Such a comparison is given in Fig. 7(a) and (b), which shows rollout predictions for streamwise and vertical velocity components respectively for three different Reynolds numbers. The rollouts are sampled at three fixed spatial (probe) locations in the step cavity region, where the BFS dynamics are most complex: probe 1 is the closest to the step anchor, and reflects the impact of flow separation on the unsteady dynamics, whereas probes 2 and 3 are successively further downstream.
@@
The rollouts ultimately confirm predictive consistency between the interpretable GNN and the corresponding baseline – in almost all cases, the predictions provided by the interpretable model very closely resemble those of the baseline. Interestingly, the rollouts in Fig. 7 reveal noticeable error buildup in the near-step regions as opposed to further downstream. The degree of error buildup is most prominent at the highest tested Reynolds number, where the onset of instability starts at about 20 rollout steps; this is consistent with the trends in Fig. 6, which implies that accumulation of single-step error is most pronounced at the highest Reynolds numbers.
@@
While the rollout accuracy, particularly near the step, is less than ideal, two important points must be stressed: (a) the GNN timestep $\Delta t$ is prescribed here to be significantly larger than the CFD timestep (as discussed in the previous section); and (b) error accumulation is expected since the objective function used during training and interpretability enhancement only accounts for the single-step error. Although inclusion of multiple rollout steps during training has been shown in recent work to significantly improve GNN prediction horizons for baseline models [6], such investigations into long-term stability improvements were not explored in this work, as they significantly increase training times and require complex application-specific training schedules. Since the scope of this work is interpretability enhancement with respect to a baseline, the objective of Fig. 7 is to highlight how appending the interpretable sub-sampling module to the baseline GNN does not incur significant accuracy penalties in both single-step and rollout contexts. As a result, it can be concluded that the added benefit of masked field interpretability is a meaningful addition to the black-box model.
@@
Figure 7: (a) Rollout predictions for streamwise velocity component at fixed probe locations for baseline GNN (blue squares) and interpretable GNN (RF=16, red circles). Solid black line indicates target value. Rows correspond to different initial conditions sourced from the indicated Reynolds numbers; columns correspond to different probe locations at specified x-locations away from step anchor (y-locations of probes coincide with the step height). (b) Same as (a), but for vertical velocity component. (c) Masked field after 1 rollout step for Re=45,589. (d) Masked field after 16 rollout steps for Re=45,589; inlay shows zoom-in on a portion of the masked field.
@@
Figure 7(c) and (d) provide masked field visualizations from the Re=$45,589$ predictions produced by the interpretable GNNs, and show the effect of rollout error accumulation on the masked field structures. More specifically, the masked field at the first rollout step in Fig. 7(c) isolates largely coherent structures in the step cavity, as observed in the previous examples (see Fig. 5), whereas the masked field after 16 rollout steps (Fig. 7(d)) includes the emergence of high frequency checkerboard-like patterns that are largely non-physical. These visualizations show a direct correspondence between rollout errors and spatial coherency observed in the identified masked field structures – although more investigation to this end is required, the advantage is that the masked field is generated directly by the model during inference, and it is possible that such high-frequency artifacts can be used to signal oncoming model error accumulation in an a-posteriori context.
@@ 3.2 Error-Identifying GNNs
The results above showcased the interpretability enhancement provided by the interpretable model from the angle of feature extraction. More specifically, the structures identified by the interpretable GNNs through the masked fields serve as an accessible link between the model architecture and the optimization goal (the mesh-based flow forecast).
The goal of this section is to build on the above by adding more capability to the identified masked fields from the modeling perspective (i.e., providing utility beyond feature extraction). This is accomplished by adding a regularization term, representing a mean-squared error budget, to the objective function during the interpretability enhancement process. It is shown below how the inclusion of this relatively simple regularization forces the identified nodes in the masked field to not only produce coherent features, but also to coincide with regions that contribute most significantly to the interpretable GNN forecasting error, a novel capability not guaranteed in the standard interpretability enhancement procedure.
@@ 3.2.1 Error Identification via Regularization
A novel error identification strategy is accomplished using a budget regularization that modifies the objective function used during the interpretability enhancement procedure. This modified objective function adds a regularization term to the standard single-step prediction objective of Eq. 3, and is given by
@@
Figure 8: (Left) Average masked field budgets (Eq. 6) versus testing Reynolds number produced by interpretable GNNs using increasing values of $\lambda$. Black curves correspond to streamwise velocity component, and red to vertical velocity component. (Right) Normalized RMSE (Eq. 4) versus testing Reynolds number for baseline GNN (black) and interpretable GNNs trained with different levels of budget regularization scaling factor $\lambda$ (markers). All interpretable GNNs in figure use $RF=16$. RMSE curves shown for vertical velocity component (trends are consistent for streamwise component).
In Eq. 6, the quantity ${\cal B}$ is the masked field budget, and represents the contribution to the total mean-squared error for the single-step forecast from the nodes identified in the masked field ${\bf M}_{m}^{i}$, where the subscript $m$ denotes dependence on the input flowfield at the $m$-th time step (${\bf X}_{m}$), and the superscript $i$ indexes a node in the graph. As per the definition of the masked field, ${\bf M}_{m}^{i}$ is 1 if the node is identified through the Top-K pooling procedure, and 0 otherwise. Since the regularization term ${\cal L}_{B}$ is cast as the inverse of the budget ${\cal B}$, a model with a small value of ${\cal L}_{B}$ indicates that a large amount of forecasting error is contained in the nodes identified through the masked field. The parameter $\lambda$ in Eq. 5 dictates the influence of the regularization term during the interpretability enhancement stage, and therefore plays a critical role. Note that setting $\lambda=0$ recovers the non-regularized interpretability enhancement strategy discussed throughout Sec. 3.1. Through the dependence of ${\cal B}$ on the masked field, the objective is to ensure that – if the regularization strategy succeeds in reducing ${\cal L}_{B}$ – the nodes responsible for a majority of the forecasting error are accessible as an explicit function of the input mesh-based flowfield ${\bf X}_{m}$, resulting in a type of a-posteriori error tagging strategy.
@@ 3.2.2 Impact of Regularization
The effect of the scale factor $\lambda$ on the budget-regularized interpretability enhancement is shown in Fig. 8 (left), which plots average budget values (single-step prediction error contained in the masked fields) conditioned on both testing set Reynolds number and velocity component for various interpretable GNN configurations. The figure illustrates (a) how various degrees of error tagging capability are indeed being provided through the budget regularization procedure in direct proportion to $\lambda$, and (b) the expected relationship between the node reduction factor and budget (higher $RF$ gives lower budget contributions). For example, the standard interpretability enhancement strategy corresponding to $\lambda=0$ produces masked fields that describe a noticeably lower amount of model error than the regularized ($\lambda>0$) counterparts. Interestingly, the implication is that without budget regularization, the generated masked fields – even for a relatively large node reduction factor of $RF=4$, meaning that $25\%$ of the total nodes are sampled – account for only roughly half of the forecasting error contribution. On the other hand, as $\lambda$ is progressively increased from $10^{-4}$ to $10^{-2}$, the budget values increase across the board. The $\lambda=10^{-2}$ case nearly reaches a saturation point, where a majority of the single-step forecasting error in most cases is described by the identified nodes. The budget curves are almost flat, with some exceptions at the higher tested Reynolds numbers, implying that the regularization strategy learns a robust and interpretable error-identifying mechanism.
@@
Although the increase in $\lambda$ translates to an increase in error-identifying capability in the interpretable GNNs, Fig. 8 (right) shows how the regularization incurs an accuracy tradeoff. More specifically, Fig. 8 (right) shows effects of the $\lambda$ parameter on the interpretable GNNs in terms of single-step normalized RMSE values (Eq. 4) conditioned on the same testing set Reynolds numbers. When the budget regularization is introduced, the scaling parameter effectively increases the forecasting errors for all observed Reynolds numbers, implying a type of accuracy tradeoff with respect to the additional modeling capability provided by the budget regularization. This is overall consistent with the $\lambda$-scaling process, in that a greater emphasis on minimizing the inverse budget via high $\lambda$ necessitates less optimization focus on the direct forecasting error objective. This attribute is directly apparent in Fig. 8 (right) – as $\lambda\rightarrow 0$, the RMSE curves converge to the standard interpretable GNN, which achieves lower forecasting error relative to the baseline GNN used to initiate the interpretability enhancement procedure.
@@
To further assess the effect of budget regularization in a more practical inference-based setting, Fig. 9 displays forecasting results in terms of masked field budget and predicted velocity versus time step using an initial condition sourced from the Re=35,392 testing set trajectory. The figure compares interpretable GNN forecasts in single-step and rollout prediction configurations, with models corresponding to node reduction factors (RFs) of 16 (the adaptive pooling procedure isolates sub-graphs comprised of 16x fewer nodes than the original).
@@
The single-step forecasts largely mimic the trends observed in Fig. 8 (left), in that the regularization imposed during interpretability enhancement manifests through an increase in error budget during inference with higher $\lambda$. These single-step forecasts, which assume no accumulation of model forecasting error through forward time steps, highlight the control provided by the regularization procedure by the $\lambda$ parameter. More specifically, Fig. 9 indicates diminishing returns in the regularization procedure with respect to $\lambda$: for example, the jump in budget when moving from $\lambda=0$ to $\lambda=10^{-4}$ is on average higher than when moving from $\lambda=10^{-3}$ to $\lambda=10^{-2}$. The rollout trends, which do account for the accumulation of single-step forecasting error through trajectory generation, further point to this aspect of diminishing returns, and directly reveal how the added emphasis on budget regularization through higher $\lambda$ values incurs an accuracy tradeoff. For example, the interpretable GNN trained with $\lambda=10^{-2}$ departs from the target solution much earlier than the others, which is consistent with the fact that it places more emphasis on minimizing the inverse budget during interpretability enhancement – ultimately, the cost of having a more “useful” masked field through the ability to tag regions of high error is a drop in forecasting accuracy and stability. Interestingly, although the correlation between budget and the $\lambda$ parameter is present for the first few rollout steps, this correlation decays once error starts to pollute the predictions. In other words, the budgets in the regularized models eventually tend towards the budgets produced by the non-regularized model. Since the baseline and interpretability enhancement procedures utilize single-step training objectives, the impact of error accumulation after a single forecasting step is high – as mentioned earlier, these effects can be mitigated with more robust (albiet more expensive) training strategies that account for the error of more than one future time-step in the objective function.
@@
Figure 9: (Top row) Masked field budgets (Eq. 6) versus time step for single-step (left plot) and rollout (right plot) predictions using initial condition from Re=35,392 testing set trajectory. interpretable GNNs have RF=16. (Bottom row) Corresponding velocity field predictions (vertical component) sampled at a single spatial location (probe 2, same as used in middle row of Fig. 7(b)). The $\lambda=-L_{\text{BL}}$ setting (green squares) is discussed in Sec. 3.2.3.
@@
Figure 10: (Left) Masked fields produced by interpretable GNNs (RF=16) of increasing $\lambda$ settings for a given input ${\bf X}_{m}$ (input comes from trajectory described in Fig. 9). (Middle) Streamwise component of the forecast error (${\bf X}_{m+1}-\widetilde{\bf X}_{m+1}$) after a single step prediction for each of the corresponding interpretable models. (Right) Vertical component of the error.
@@
It should be noted that for small values of $\lambda$ (e.g., $\lambda=10^{-4}$ in Fig. 9), the regularization in Eq. 5 produces error-identifying masked fields at relatively minimal expense in terms of model accuracy with respect to the baseline GNN, thereby offering a unique interpretable modeling advantage from the perspective of error tagging. This advantage is more apparent when considering Fig. 10, which shows masked fields produced by the interpretable GNNs used to make the predictions in Fig. 9. The introduction of $\lambda$ incurs a noticeable change to the identified structures in the masked field; more specifically, the non-regularized $\lambda=0$ interpretable GNNs isolate structures in the far downstream region, whereas the regularized models isolate more nodes in the near-step vicinity, particularly near the corner at which flow separation is initiated. The difference in masked field structures emphasizes the difference in interpretable GNN objectives, in that the identified nodes in the $\lambda>0$ models represent regions in which the predicted mesh-based flowfield $\widetilde{X}_{m+1}$ will contain a majority of the model error. This is evidenced in the corresponding error fields, which showcases the effect of budget regularization on the localization of forecasting error towards the masked regions.
@@
Additionally, the structures in the masked fields from $\lambda=10^{-4}$ to $\lambda=10^{-2}$ in Fig. 10 are largely similar, which allows the user to tune the degree of error identification capability provided by the interpretable GNN. For example, the $\lambda=10^{-4}$ model offers (a) comparable performance to the baseline GNN, (b) interpretability through the identification of coherent structures in the masked field, a utility also provided in the non-regularized $\lambda=0$ model, and (c) the ability to attribute the identified regions to expected sources of forecasting error. Although this latter capability is limited to the scope of the prediction horizon of the model (this is apparent through the rollout results in Fig. 9), the inclusion of the budget-based regularization reveals a promising pathway for understanding and isolating sources of model error during inference through the adaptive pooling strategy.